# Quasar Pairs as Large–Scale Structure tracers

### v1

@@ 5 Cross-correlation Quasar-Pair
We propose an additional method to confirm that the observed correlation is physically strong. It involves partitioning the original subsample within the redshift window $z\in[2.2,2.8]$ into two distinct datasets, which are created by applying a cutoff on the logarithm of the sum of the luminosities. By adopting the aforesaid approach, we can further assess whether our results maintain physical relevance across different partitions. In fact, Fig. (15) shows that the cross-correlation signal persists regardless of the cut-off or partitioning method employed. Although the error bars may appear larger, even with the same bootstrap method applied, the larger size is primarily associated with the reduced sample size after partitioning based on the following specified criterion, $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}\lessgtr 14.548$. Such observation underscores the resilience of the correlation between different approaches, reinforcing the validity of our findings despite the variations in data partitioning. As anticipated, linear regression may be less reliable in certain scenarios. For the case where $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}>14.548$, the slope is found to be $\gamma=(2.00\pm 0.40)$ and the correlation length is $s_{0}=(16.00\pm 1.70)$, with a coefficient of determination $R^{2}\simeq 0.80$. In contrast, for the opposite case where $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}\leq 14.548$, the estimations yield $\gamma=(2.00\pm 0.63)$, $s_{0}=(16.00\pm 4.15)$, and $R^{2}\simeq 0.68$.
Before concluding this section, we investigate the cross-correlation function within the redshift range of $1.2\leq z\leq 2.8$. Our objective is to evaluate the persistence of the cross-correlation signal. We found a significant increase in the number of quasar pairs, with $N_{p}=2,777$, as a result of employing the section method outlined in Sect. 4. Figure (16) illustrates the cross-correlation function within the specified interval, juxtaposing its signal against the autocorrelation of quasars and the autocorrelation of isolated quasars, those that do not belong to any pair. We also verified that the cross-correlation between the pairs and the isolated quasars does not yield an amplified signal compared to the full cross-correlation between the complete sample within this interval and the detected pairs. In summary, we found that the signal of the cross-correlation between pairs and quasars consistently exceeds the signal of the autocorrelation of quasars. This observation will be further investigated in the next section by an examination of the convergence CMB lensing signal. Moreover, we will explore whether the amplitude of the convergence signal is mainly attributable to the more luminous quasar pairs or to the less luminous ones.
We conclude the section by addressing two observational and numerical points pertinent to the detection of quasar pairs. Firstly, we pose a fundamental question: Is the SDSS camera capable of effectively resolving two quasars that constitute a confirmed pair? This query is essential for understanding the limitations of observational data and the ability of the SDSS to distinguish quasar pairs, which is crucial for accurate pair identification and subsequent analysis. To answer that question, we consider a conservative angular resolution for the SDSS telescope Lyke (2020), with $\theta\geq\theta_{SDSS}=5\text{arc sec}$ , and a mean redshift value of $z_{\text{mean}}=2.4$. The physical distance corresponding to the SDSS angular resolution yields approximately $d\simeq 0.48\text{Mpc}$. Such findings are significant in light of previous observations, which indicated that the distribution of distances between quasar pairs is around $d=1.05\text{Mpc}$. Thus, we can conclude that the SDSS camera is capable of resolving these objects.
A second numerical estimation focuses on the comoving numerical density of quasars and quasar pairs within the redshift interval $[2.2,2.8]$. The fraction of the sky surveyed is approximately $0.65\%$, leading to a numerical (comoving) density of $n_{q}\simeq 8.8\times 10^{-9}\text{Mpc}^{-3}$. This value aligns well with the findings reported in Croom (2005); namely, $n_{q}\simeq 8.2\times 10^{-9}\text{Mpc}^{-3}$. Taking into consideration a minor sky fraction for the quasar pairs $f_{\text{sky, pair}}=0.54\%$, we observe a notable reduction in the estimated numerical (comoving) density of these pairs. It approximately yields $n_{\text{pair}}\simeq 3.2\times 10^{-11}\text{Mpc}^{-3}$. The latter finding highlights the significant influence that observing a restricted area of the sky, combined with a small sample of detected pairs, has on the estimated density of quasar pairs, particularly in contrast to the densities associated with individual quasars Croom (2005).
@@ 6 CMB Lensing Signal
We conduct an analysis of the CMB convergence signal across the redshift interval $z\in[1.2,2.8]$ (associated with the full sample) and estimate the cosmic variance within this range [see Fig. (17)]. The signal appears to fall within the bounds of the cosmic variance estimation, however the quasar pairs in the redshift range of $[1.2,2.2]$ exhibit a larger amplitude compared to the overall sample and the cosmic variance; confirming the previous analysis on cross-correlation function quasar-quasar pairs [cf. Fig. (16)]. This behavior is congruent with the maximum amplitude expected from the theory, taking into account that the source is located at $z\sim 1100$. Therefore, we further study the redshift range $z\in[1.2,2.2]$ where $2126$ pairs were established in order to maximize the signal-to-noise ratio.
@@
Figure 18: Top panel: Radial Convergence Profiles for the quasar sample in the redshift range $1.2\leq z\leq 2.2$ dividing into pairs of quasars, high luminosity and low luminosity isolated quasars. Bottom Panel: Radial Convergence Profiles for the pairs of quasars dividing into high luminosity quasar pairs and low luminosity quasar pairs. In both panels, the division is done using the median of $L_{1}+L_{2}$ for each sample.
@@ 7 Summary and Discussions
In this study, we construct a sub-sample from the SDSS DR16 quasar catalog, focusing on a redshift range ($z\in[2.2,2.8]$) where the sample exhibits a good completeness factor. We conducted a comprehensive analysis of this subsample, with particular emphasis on the detection of quasar pairs. To achieve this, we first calculated the projected distance in Mpc between the two members of each quasar pair. We performed an extensive exploratory analysis of the luminosity distribution of the quasar pairs, examining how they distinguish themselves from both fake and simulated pairs. Subsequently, we computed the cross-correlation function between the quasar pairs and the overall quasar population, revealing a distinctly amplified signal above the background of isolated quasars. We investigated how this correlation is influenced by various scenarios, including partitioning the redshift interval into two segments, separating the data according to the radio luminosity ratio $L_{1}/L_{2}$, and applying a cutoff in $\log_{10}(L_{1}+L_{2})/L_{\odot}$. The main amplitude of the cross-correlation signal remained consistent under these conditions. We also addressed the observational capabilities for detecting quasar pairs using SDSS and contrasted the numerical comoving density of the quasar pairs with that of the quasar subsample.
We expanded upon previous studies by considering a lower redshift window ($1.2\leq z\leq 2.2$), achieving a higher detection rate of quasar pairs ($N_{p}=2126$). As a result of this analysis, we confirmed that the amplitude of the cross-correlation between pairs and quasars consistently surpasses the signal of the auto-correlation of quasars [cf. Fig. 16]. Subsequently, we validated this detection through an additional observational test, concentrating on the amplitude of the CMB lensing signal counterpart by examining the convergence $\kappa$ signal. We found that the most luminous pairs of quasars are the primary contributors to the total signal for the sample [cf. Fig. 18]. In addition, the CMB convergence signal appears to be within the limits set by the cosmic variance estimation across the redshift interval $z\in[1.2,2.8]$. However, quasar pairs located within the $z\in[1.2,2.2]$ range show a higher amplitude relative to both the total sample and the cosmic variance [see Fig. (17)].
@@ Acknowledgements.
We acknowledge the use of the SDSS DR16 catalog https://www.sdss4.org/dr16/algorithms/qso_catalog, Lyke (2020). We used several packages such as Astropy Robitaille (2013), Corrfunc Sinha (2023), Numpy Oliphant (2006), Matplotlib Hunter (2007), Scipy Virtanen (2020), Healpy Zonca (2019), HEALPix Górski et al (2005), Pandas McKinney (2010). The authors thank CNPq, FAPES, and Fundação Araucária for financial support.

### v2

@@ 6 Cross-correlation Quasar-Pair
We suggest an alternative approach to validate that the observed cross-correlation is physically significant and not merely a consequence of the chosen redshift bins. It involves partitioning the original subsample within the redshift window $z\in[2.2,2.8]$ into two distinct datasets, which are created by applying a cutoff on the logarithm of the sum of the luminosities. By adopting the aforesaid approach, we can further assess whether our results maintain physical relevance across different partitions bu using the luminosities of the quasar pairs as an input. In fact, Fig. (16) shows that the cross-correlation signal persists regardless of the cut-off or partitioning method employed. Although the error bars may appear larger, even with the same bootstrap method applied, the larger size is primarily associated with the reduced sample size after partitioning based on the following specified criterion, $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}\lessgtr 14.548$. Such observation underscores the resilience of the correlation between different approaches, reinforcing the validity of our findings despite the variations in data partitioning. As anticipated, linear regression may be less reliable in certain scenarios. For the case where $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}>14.548$, the slope is found to be $\gamma=(2.00\pm 0.40)$ and the correlation length is $s_{0}=(16.00\pm 1.70)$, with a coefficient of determination $R^{2}\simeq 0.80$. In contrast, for the opposite case where $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}\leq 14.548$, the estimations yield $\gamma=(2.00\pm 0.63)$, $s_{0}=(16.00\pm 4.15)$, and $R^{2}\simeq 0.68$.
These findings warrant further exploration, which will be undertaken in the next section through an investigation of the convergence CMB lensing signal 7. Additionally, we will assess whether the amplitude of the convergence signal is primarily influenced by the more luminous quasar pairs or the less luminous ones.
@@ 7 CMB Lensing Signal
We conduct a general analysis of the CMB convergence signal across the redshift interval $z\in[1.2,2.8]$ (associated with the full sample) and estimate the cosmic variance within this range [see Fig. (17)]. The signal appears to fall within the bounds of the cosmic variance estimation, however the quasar pairs in the redshift range of $[1.2,2.2]$ exhibit a larger amplitude compared to the overall sample and the cosmic variance; confirming the previous analysis on cross-correlation function quasar-quasar pairs [cf. Fig. (13)]. This behavior is congruent with the maximum amplitude expected from the theory, taking into account that the source is located at $z\sim 1100$. Therefore, we further study the redshift range $z\in[1.2,2.2]$ where $2126$ pairs were established in order to maximize the signal-to-noise ratio.
@@
Figure 18: The radial convergence profiles for quasar pairs, categorized by luminosity levels (dotted lines for high and low luminosity), are shown in the redshift range $1.2\leq z\leq 2.2$. For comparison, the profiles of isolated quasars are shown with dashed lines. The results indicate that the amplitude of the radial convergence for quasar pairs consistently exceeds that of isolated quasars. Moreover, the high luminosity quasar pairs appear to contribute significantly to a signal that surpasses the expected cosmic variance
@@ 8 Summary and Discussions
In this study, we construct a sub-sample from the SDSS DR16 quasar catalog, focusing on a redshift range ($z\in[1.2,2.8]$) where the sample exhibits a good completeness factor. We conducted a comprehensive analysis of this subsample, with particular emphasis on the detection of quasar pairs. To achieve this, we first calculated the projected distance in Mpc between the two members of each quasar pair.
We performed an extensive exploratory analysis of the luminosity distribution of the quasar pairs, examining how they distinguish themselves from both control and simulated pairs. Moreover, we computed the cross-correlation function between the quasar pairs and the overall quasar population, revealing a distinctly amplified signal above the background of isolated quasars. We investigated how this correlation is influenced by various scenarios, including partitioning the redshift interval into two segments, separating the data according to the radio luminosity ratio $L_{1}/L_{2}$, and applying a cutoff in $\log_{10}(L_{1}+L_{2})/L_{\odot}$. The main amplitude of the cross-correlation signal remained consistent under these conditions. We also addressed the observational capabilities for detecting quasar pairs using the SDSS camera, discussed the possibility of contamination of the quasar pair sample by foreground galaxies, and contrasted the numerical comoving density of the quasar pairs with that of the quasar subsample.
We have demonstrated that the CMB convergence signal remains within the bounds defined by cosmic variance estimates for the redshift interval $z\in[1.2,2.8]$. When we refine our focus to a narrower redshift range, $z\in[1.2,2.2]$, the CMB convergence signal associated with quasar pairs exhibits a heightened amplitude, surpassing not only the total quasar pair sample but also the estimated cosmic variance, as illustrated in Fig. (17). Furthermore, our analysis confirms that the amplitude of the cross-correlation between quasar pairs and individual quasars consistently exceeds that of the quasar auto-correlation within the same redshift range (see Fig. 13). By integrating the cross-correlation function with the CMB convergence signal, we achieved a substantially enhanced detection rate of quasar pairs at lower redshifts, culminating in a total of $N_{p}=2126$.
We further validated our detection through an additional observational test, focusing on the amplitude of the CMB lensing signal associated with quasar pairs distinguished by high and low luminosities, using the mean luminosity as the cutoff between these two groups. Our analysis revealed that the most luminous pairs of quasars are the dominant contributors to the overall signal for the sample, as illustrated in Fig. 18.
@@ Acknowledgements.
We acknowledge the use of the SDSS DR16 catalog https://www.sdss4.org/dr16/algorithms/qso_catalog, Lyke et al. (2020). We used several packages such as Astropy Robitaille et al. (2013), Corrfunc Sinha (2023), Numpy Oliphant (2006), Matplotlib Hunter (2007), Scipy Virtanen et al. (2020), Healpy Zonca et al. (2019), HEALPix Górski et al. (2005), Pandas McKinney (2010). The authors thank CNPq, FAPES, and Fundação Araucária for financial support. The authors would like to express their gratitude for the valuable comments provided by the referee, which significantly contributed to enhancing the quality and clarity of the paper.